Sarah invested a certain amount of money into a savings account, and invested 3 times that amount into a mutual fund. If after 1 year the savings account earned 5% interest and the mutual fund earned 11% interest, the total interest earned from the two investments was $152. How much did Sarah initially invest in each fund???
Answer by ptaylor(2198)

Let x=amount that Sarah invested in the savings account
Then 3x=amount that she invested in the mutual fund
Interest(I)=Principal(P) times Rate(R) times Time(T) or I=PRT
(1) Interest on savings account=x*(0.05)*(1)=0.05x
(2) Interest on mutual fund=3x*(0.11)*(1)=0.33x
Now we are told that (1)+(2)=$152. So our equation to solve is:
0.05x+0.33x=152 combine like terms
0.38x=152 divide both sides by 0.38
x=$400-------------amount that Sarah invested in the savings account
3x=3*$400=$1200----------amount that she invested in the mutual fund
CK
0.05*400+0.11*1200=$152
20+132=152
152=152

Hope this helps---ptaylor